OpenStudy (anonymous):

Which of the following is one of the factors of 2x2 + 2x – 144?

OpenStudy (anonymous):

these are the options: 2 x + 9 All of the above x – 8

OpenStudy (anonymous):

2x2 + 2x – 144 = 2[x^2 + x - 72] = 2[x^2 + 9x - 8x - 72] = 2[x(x + 9) -8(x + 9)] = 2[(x - 8)(x + 9)] = 2 (x - 8)(x + 9) from this it is evident that all are factors of the given quadratic polynomial so ans is "All of the above"